I have a problem for ya
Solve for the value of x: (4x-35)/3 = 9(1-x)

Answer by Nate(3500)

(4x - 35)/3 = 9(1 - x)
4x - 35 = 27(1 - x) mult. by 3
4x - 35 = 27 - 27x
31x = 62
x = 2